2. A 5-ft square, 2-ft deep spread footing is subjected to a concentric vertical load of 60 k and an overturning moment of 30 ft-k. The overturning moment acts parallel to one of the sides of the footing, the top of the footing is flush with the ground surface, and the ground water table is at a depth of 20 ft below the ground surface. Compute the minimum and maximum bearing pressures, d show the distribution of bearing pressure in a sketch.
media%2F6ef%2F6ef590de-aa09-4d24-81e2-f0
Answer #1

Step 1:

Given:

M=30 ft-kips

P=60 kips

Step 2:

Calculate the ecentricity

e=M/P

e=30/60

e=0.50 ft

Step 3:

Check for the pore water pressure

As the water is at a great depth it woint have any effect on the bearing pressure

Step 4:

Calculate the minimum and maximum pressures:

qmax=P/B2(1+6e/b)=30/52(1+0.50/5)=1.32 kips/ft2

qmin=P/B2(1-6e/B)=30/52(1-0.50/5)=1.08 kips/ft2